OpenStudy (anonymous):

Amy has deposited $661 in a savings account that earns interest at a rate of 3.1% compounded monthly. What will the account balance be in 15 years?

OpenStudy (wolf1728):

First we need the formula to convert monthly interest to annual interest. Annual interest = (1 + .031/12)^12 Annual interest = (1.0025833333)^12 Annual interest = 3.1444273306 % Okay that's the first step.

OpenStudy (anonymous):

use the annual compounding equation \[FV = PV*(1+i)^{n}\]

OpenStudy (anonymous):

FV=661(1+3.144)^15

OpenStudy (wolf1728):

Well seems Shorthell posted the next step.

OpenStudy (anonymous):

so then it would be 661(4.144)^15

OpenStudy (anonymous):

when you plug in 'i' remember that a percent is actually a decimal, so 3.1% is actually 0.031

OpenStudy (wolf1728):

NO - INTEREST RATE IS INCORRECTLY ENTERED

OpenStudy (wolf1728):

Ugh hit the caps key - sorry

OpenStudy (anonymous):

661(1+0.031)^15

OpenStudy (anonymous):

661(1.031)^15

OpenStudy (anonymous):

661(1.58)

OpenStudy (wolf1728):

FV=661*(1+.03144)^15 That looks correct

OpenStudy (anonymous):

@Shorthell you're doing it right

OpenStudy (anonymous):

then 661(1.58)=1044.38

OpenStudy (anonymous):

then add 661 to 1044.38

OpenStudy (anonymous):

i must have done something wrong because its not one of my answers

OpenStudy (wolf1728):

I get 1,051.63 I think that decimal rounded to 1.58 makes the difference

OpenStudy (anonymous):

oh, the 1044.38 is the "future value" that 661 grows into, so it's just 1044.38

OpenStudy (anonymous):

i think wolf is right i did round though and got a complete different answer.

OpenStudy (wolf1728):

thanks Shorthell

OpenStudy (anonymous):

no prob but the closest answer is yours but is off by a lil the answer is 1051.69

OpenStudy (anonymous):

close enugh for me! ^_^ good job @Shorthell

OpenStudy (wolf1728):

Okay I recalculated. I used the 3.144 interest but the actual interest is 3.1444273306 .03144 ^ 15 = 1.5909610408 .031444273306 ^ 15 = 1.5910599150921 see how "dramatically" that changes the total?

OpenStudy (anonymous):

what would the final answer be?

OpenStudy (wolf1728):

The final answer is $1,051.69

OpenStudy (anonymous):

I thought so haha

OpenStudy (wolf1728):

I'm glad you came back to check. Yes, I just had to know where that difference was.